# Structured Spectral Reasoning for Frequency-Adaptive Multimodal Recommendation

### v2

@@ Abstract
Multimodal recommendation aims to integrate collaborative signals with heterogeneous content such as visual and textual information, but remains challenged by modality-specific noise, semantic inconsistency, and unstable propagation over user–item graphs. These issues are often exacerbated by naive fusion or shallow modeling strategies, leading to degraded generalization and poor robustness. While recent work has explored the frequency domain as a lens to separate stable from noisy signals, most methods rely on static filtering or reweighting, lacking the ability to reason over spectral structure or adapt to modality-specific reliability. To address these challenges, we propose a Structured Spectral Reasoning (SSR) framework for frequency-aware multimodal recommendation. Our method follows a four-stage pipeline: (i) Decompose graph-based multimodal signals into spectral bands via graph-guided transformations to isolate semantic granularity; (ii) Modulate band-level reliability with spectral band masking, a training-time masking with prediction-consistency objective that suppresses brittle frequency components; (iii) Fuse complementary frequency cues using hyperspectral reasoning with low-rank cross-band interaction; and (iv) Align modality-specific spectral features via contrastive regularization to promote semantic and structural consistency. Experiments on three real-world benchmarks show consistent gains over strong baselines, particularly under sparse and cold-start settings. Additional analyses indicate that structured spectral modeling improves robustness and provides clearer diagnostics of how different bands contribute to performance.
@@ 1 Introduction
Multimodal recommendation liu2024rec ; malitesta2024formalizing ; xu2024mmpoi ; xu2025mentor has become a central paradigm in modern information systems. It leverages auxiliary signals such as visual content, textual metadata and structural interactions to improve user preference modeling and content understanding he2021click ; zhou2023comprehensive ; yang2023based ; wei2023multi ; wu2024towards . These signals provide complementary semantics but also introduce modality-specific noise liu2024multimodal ; xv2024improving ; qi2025seeing , semantic misalignment yi2025enhancing ; luo2024molar and representational redundancy ong2024spectrum ; yang2023multimodal ; luo2024qarm . The difficulties become more pronounced in graph-based architectures where user–item relations are propagated through neighborhoods wang2020disentangled ; liu2023multimodal ; liu2023megcf . In this setting, spurious cues from particular modalities, for example stylistic image patterns or ambiguous wording, can diffuse across the graph and affect otherwise unrelated nodes ye2024mm ; luo2024survey ; xu2023causal ; fang2024counterfactual . The result is unstable representations and degraded performance, especially under cold-start or noisy conditions. Many existing approaches fuse modalities in the raw feature space wang2023cl4ctr ; li2024multimodal ; kim2024self , which limits control over how different semantic granularities interact across the graph. These challenges call for a structured approach to semantic disentanglement that isolates stable signals while suppressing unstable ones kim2025disentangling ; lin2025contrastive ; su2025modality ; zhou2024disentangled .
@@
Recently, frequency-domain modeling has gained traction as a promising perspective xia2025frequency ; peng2022less ; fan2024hierarchical . By transforming signals into spectral components, models can expose structural patterns across different frequency bands luo2024dual ; zhang2023contrastive ; shin2024attentive . Low-frequency components often encode smooth, global semantics such as long-term user intent or item popularity, while high-frequency components capture fine-grained variations or local inconsistencies du2023frequency ; huang2021sliding ; han2024efficient . This decomposition enables models to analyze and potentially filter modality-specific signals with greater precision. However, advanced approaches such as SMORE ong2024spectrum and ChebyCF kim2025graph largely rely on static reweighting or low-pass filtering and therefore lack explicit reasoning about band semantics or cross-frequency dependencies. Moreover, the use of frequency decomposition itself is often heuristic, without a principled mechanism to separate informative bands from noisy ones. Consequently, useful detail may be suppressed and unstable components may be retained, particularly when semantics vary across items and users.
@@
These limitations are particularly acute in multimodal graph recommendation, where heterogeneous modalities interact within a topologically entangled space xu2024sequence ; li2024align ; yi2024unified ; bai2024multimodality . Visual and textual semantics differ in granularity and abstraction, and they also differ in spectral distribution, which makes naïve fusion or static attention insufficient ong2024spectrum ; yu2023xsimgcl ; liu2024alignrec . In this setting, robust representations should support adaptive modulation across frequency bands so that components influence one another based on semantic content and contextual relevance du2023frequency ; han2024efficient ; du2023contrastive . They should also enforce band-wise consistency across modalities to prevent collapse or redundancy within the decomposed space wang2024decoupling ; swetha2023preserving . However, current architectures lack a unified mechanism to model these structured relationships across frequency, modality and graph topology. At the same time, any solution must remain computationally tractable on large graphs and compatible with standard training pipelines. It should also provide clear diagnostics at the band level so that practitioners can inspect which frequencies and modalities drive recommendations. This motivates the need for approaches that account for band-level structure and maintain robustness across modalities.
@@
To address these challenges, we propose Structured Spectral Reasoning (SSR) for frequency-aware multimodal graph recommendation. SSR follows a four-stage pipeline that decomposes, modulates, fuses and aligns signals within a shared spectral coordinate system. Motivated by the observation that frequency mapping reveals band-wise structure capturing collaborative semantics and modality-specific detail, we organize learning around band-level operations and supervision. Instead of treating bands as static or lightly reweighted features, SSR converts inputs into spectral representations and then reasons about their informativeness and stability. We introduce Spectral Band Masking (SBM), a training-time band-level perturbation with a prediction-consistency objective that reduces reliance on brittle bands and improves robustness. We also design a graph-compatible hyperspectral operator (G-HSNO) that models cross-band and cross-modality dependencies through a compact low-rank parameterization. In addition, a spectral contrastive objective promotes band-wise cross-modal consistency without inference overhead. Together, these components provide a cohesive and efficient approach to structured spectral modeling for multimodal graphs.
@@ 2.1 Multi-modal Recommendation
Multimodal recommendation integrates auxiliary content such as images, text and audio to strengthen user preference modeling, especially under sparse supervision he2016vbpr ; chen2019personalized ; yang2024variational . Early work augments collaborative filtering with side content deldjoo2021content ; xu2023musenet ; kang2017visually . Subsequent methods build modality-aware graph architectures to propagate and fuse signals along user–item relations wei2019mmgcn ; wei2020graph ; wang2021dualgnn ; yu2022graph ; han2022modality ; tao2022self . To mitigate cross-modal misalignment and semantic noise, recent approaches employ contrastive learning zhou2023bootstrap ; yi2022multi , regularization zhang2021mining ; zhou2023tale and intention-aware designs yang2024multimodal ; guo2022topicvae . AlignRec liu2024alignrec enforces modality–ID consistency, while PromptMM wei2024promptmm enables lightweight adaptation. Generative strategies, including diffusion models jiang2024diffmm , adversarial alignment liang2022semantic and large language models zhao2025hierarchical ; xia2025hierarchical ; gu2025r ; xia2025trackrec , further improve robustness under heterogeneity. In addition, empowered by the reasoning capacity of large language models chen2025tourrank ; chen2025improving and multi-agent systems wang2024macrec ; yang2025learning ; ye2024domain , recent advances such as Rec-GPT4V liu2024rec and UniMP wei2024towards leverage large vision–language models to strengthen multimodal semantic understanding and improve reasoning over heterogeneous content, showing encouraging performance in personalized recommendation.
@@ 2.2 Frequency-aware Representation Learning
Frequency-domain modeling has recently emerged as a promising alternative to time- or topology-domain learning in recommendation tasks fan2024hierarchical ; luo2024dual . In sequential modeling, various approaches leverage spectral transformations to capture user intent dynamics and mitigate temporal noise zhang2023contrastive ; du2023frequency , while others propose frequency-aware contrastive objectives or neural filtering for periodic behavior tracking yang2025farec ; zhang2025frequency . CFIT4SRec zhang2023contrastive adopts frequency-based contrastive learning frameworks, leveraging Fourier transformations to construct robust augmented views and improve representation learning. FEARec du2023frequency and HMFTRec fan2024hierarchical add spectral attention and multi-frequency modules to capture low and high frequency signals beyond augmentation. For graph-based recommendation, ChebyCF kim2025graph , SelfGNN liu2024selfgnn , and SComGNN luo2024spectral reformulate GNN propagation as spectral filtering, improving expressivity and denoising capacity. In multimodal settings, SMORE ong2024spectrum explicitly models frequency-specific modality fusion, while FREEDOM zhou2023tale and MMGCL yi2022multi implicitly introduce spectral diversity through augmentation. Nonetheless, most of these methods either treat frequency bands as static views or lack explicit modeling of band reliability and inter-band interactions. We address these gaps by pursuing adaptive, band-aware integration that regulates sensitivity to brittle components and supports interpretable cross-modal fusion.
@@ 3.1 Spectral Transformation of Multimodal Graph Signals
Multimodal recommendation integrates diverse signals such as ID embeddings and content features from images and text into a unified user–item representation. A central challenge is the semantic and statistical heterogeneity across modalities, which often live in incompatible spaces and emphasize different levels of granularity. Conventional fusion strategies, including feature concatenation and attention-based aggregation, operate in the raw feature domain and do not capture the structural relationships induced by user–item interactions. In contrast, viewing these representations as graph signals enables frequency-aware modeling in which components are disentangled by their spectral characteristics. Low-frequency components capture shared global preference patterns, whereas high-frequency components encode fine-grained, modality-specific variations. Motivated by this perspective, we transform multimodal features into a unified frequency domain where band-level alignment is feasible and fusion becomes more interpretable. This transformation provides a coherent basis to mitigate cross-modal misalignment, improve semantic compositionality, and support adaptive selection of frequency components for downstream recommendation.
Figure 1: Overall architecture of our proposed framework. The model follows a structured four-stage pipeline: (i) Decomposition performs modality-specific graph wavelet transformation to disentangle multi-frequency components; (ii) Modulation applies Spectral Band Masking (SBM) to perturb and down-weight unreliable bands in a task-adaptive manner; (iii) Fusion leverages a low-rank Graph HyperSpectral Neural Operator (G-HSNO) to reason over cross-band and cross-modal dependencies; and (iv) Alignment introduces Spectral Contrastive Regularization (SCR) to enforce semantic consistency and spectral robustness across modalities.
@@
Given a user–item bipartite graph $\mathcal{G}=(\mathcal{N},\mathcal{E})$ with adjacency matrix $\mathbf{A}$ defferrard2016convolutional ; wu2019simplifying , we use the symmetric normalized Laplacian $\mathbf{L}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$, where $\mathbf{D}$ is the degree matrix. For each node $n\in\mathcal{N}$, the associated signal (e.g., an ID embedding or a multimodal feature) is $\mathbf{x}_{n}\in\mathbb{R}^{d}$. Stacking all nodes yields $\mathbf{X}\in\mathbb{R}^{|\mathcal{N}|\times d}$. We analyze $\mathbf{X}$ in the spectral domain via the eigendecomposition $\mathbf{L}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$, where $\mathbf{U}$ contains orthonormal eigenvectors and $\boldsymbol{\Lambda}$ is diagonal with eigenvalues. The graph Fourier transform (GFT) is $\widehat{\mathbf{X}}=\mathbf{U}^{\top}\mathbf{X}$ and the inverse transform is $\mathbf{X}=\mathbf{U}\widehat{\mathbf{X}}$. Each row of $\widehat{\mathbf{X}}$ corresponds to a graph frequency (eigenmode). In the next step we group these modes into bands for downstream processing.
@@
which yields bands with comparable signal mass and avoids extreme imbalance across modes. Given the orthonormal basis $\mathbf{U}=[\mathbf{u}_{1},\ldots,\mathbf{u}_{|\mathcal{N}|}]$, the reconstruction for band $m$ is $\mathbf{X}^{(m)}=\sum_{k\in\mathcal{F}_{m}}\mathbf{u}_{k}\,\widehat{\mathbf{x}}_{k}^{\top}\,$, where $\mathbf{u}_{k}$ is the $k$-th eigenvector. Each $\mathbf{X}^{(m)}$ captures information at a specific semantic resolution, with lower-index bands concentrating smoother components and higher-index bands emphasizing localized variations. In practice, $M$ is treated as a hyperparameter. Bands may be formed using various rules, such as equal-energy quantiles of the cumulative spectrum or uniform splits in the eigenvalue domain.
@@
where $\alpha_{n}^{(m)}\in\mathbb{R}_{\geq 0}$ are band-specific weights produced from node context, for example an MLP over the ID embedding and optional band statistics, and normalized with a softmax over $m$ so that $\sum_{m}\alpha_{n}^{(m)}=1$. This allows each node to emphasize informative bands while down-weighting unstable ones. Because subsequent linear operators distribute over the sum, the fusion can be placed before or after standard graph layers without changing the rest of the pipeline.

### v3

@@ Abstract
Multimodal recommendation aims to integrate collaborative signals with heterogeneous content such as visual and textual information, but remains challenged by modality-specific noise, semantic inconsistency, and unstable propagation over user–item graphs. These issues are often exacerbated by naive fusion or shallow modeling strategies, leading to degraded generalization and poor robustness. While recent work has explored the frequency domain as a lens to separate stable from noisy signals, most methods rely on static filtering or reweighting, lacking the ability to reason over spectral structure or adapt to modality-specific reliability. To address these challenges, we propose a Structured Spectral Reasoning (SSR) framework for frequency-aware multimodal recommendation. Our method follows a four-stage pipeline: (i) Decompose graph-based multimodal signals into spectral bands via graph-guided transformations to isolate semantic granularity; (ii) Modulate band-level reliability with spectral band masking, a training-time masking with representation-consistency objective that suppresses brittle frequency components; (iii) Fuse complementary frequency cues using hyperspectral reasoning with low-rank cross-band interaction; and (iv) Align modality-specific spectral features via contrastive regularization to promote semantic and structural consistency. Experiments on three real-world benchmarks show consistent gains over strong baselines, particularly under sparse and cold-start settings. Additional analyses indicate that structured spectral modeling improves robustness and provides clearer diagnostics of how different bands contribute to performance. The code is available at https://github.com/llm-ml/SSR.git.
@@ 1 Introduction
Multimodal recommendation Liu et al. (2024c); Malitesta et al. (2024); Xu et al. (2024b, 2025) has become a central paradigm in modern information systems. It leverages auxiliary signals such as visual content, textual metadata and structural interactions to improve user preference modeling and content understanding He et al. (2021); Zhou et al. (2023a); Yang et al. (2023b); Wei et al. (2023); Wu et al. (2024). These signals provide complementary semantics but also introduce modality-specific noise Liu et al. (2024a); Xv et al. (2024); Qi et al. (2025), semantic misalignment Yi et al. (2025); Luo et al. (2024e) and representational redundancy Ong and Khong (2024); Yang et al. (2023c); Luo et al. (2024d). The difficulties become more pronounced in graph-based architectures where user–item relations are propagated through neighborhoods Wang et al. (2020); Liu et al. (2023a, b); Yang et al. (2023a). In this setting, spurious cues from particular modalities, for example stylistic image patterns or ambiguous wording, can diffuse across the graph and affect otherwise unrelated nodes Ye et al. (2024b); Luo et al. (2024b); Xu et al. (2023b); Fang et al. (2024). The result is unstable representations and degraded performance, especially under cold-start or noisy conditions. Many existing approaches fuse modalities in the raw feature space Wang et al. (2023); Li et al. (2024a); Kim et al. (2024), which limits control over how different semantic granularities interact across the graph. These challenges call for a structured approach to semantic disentanglement that isolates stable signals while suppressing unstable ones Kim et al. (2025b); Lin et al. (2025); Su et al. (2025); Zhou and Miao (2024).
@@
Recently, frequency-domain modeling has gained traction as a promising perspective Xia et al. (2025a); Peng et al. (2022); Fan et al. (2024). By transforming signals into spectral components, models can expose structural patterns across different frequency bands Luo et al. (2024c); Zhang et al. (2023); Shin et al. (2024). Low-frequency components often encode smooth, global semantics such as long-term user intent or item popularity, while high-frequency components capture fine-grained variations or local inconsistencies Du et al. (2023b); Huang et al. (2021); Han et al. (2024). This decomposition enables models to analyze and potentially filter modality-specific signals with greater precision. However, advanced approaches such as SMORE Ong and Khong (2024) and ChebyCF Kim et al. (2025a) largely rely on static reweighting or low-pass filtering and therefore lack explicit reasoning about band semantics or cross-frequency dependencies. Moreover, the use of frequency decomposition itself is often heuristic, without a principled mechanism to separate informative bands from noisy ones. Consequently, useful detail may be suppressed and unstable components may be retained, particularly when semantics vary across items and users.
@@
These limitations are particularly acute in multimodal graph recommendation, where heterogeneous modalities interact within a topologically entangled space Xu et al. (2024a); Li et al. (2024b); Yi and Ounis (2024); Bai et al. (2024). Visual and textual semantics differ in granularity and abstraction, and they also differ in spectral distribution, which makes naïve fusion or static attention insufficient Ong and Khong (2024); Yu et al. (2023); Liu et al. (2024b). In this setting, robust representations should support adaptive modulation across frequency bands so that components influence one another based on semantic content and contextual relevance Du et al. (2023b); Han et al. (2024); Du et al. (2023a). They should also enforce band-wise consistency across modalities to prevent collapse or redundancy within the decomposed space Wang et al. (2024a); Swetha et al. (2023). However, current architectures lack a unified mechanism to model these structured relationships across frequency, modality and graph topology. At the same time, any solution must remain computationally tractable on large graphs and compatible with standard training pipelines. It should also provide clear diagnostics at the band level so that practitioners can inspect which frequencies and modalities drive recommendations. This motivates the need for approaches that account for band-level structure and maintain robustness across modalities.
@@
To address these challenges, we propose Structured Spectral Reasoning (SSR) for frequency-aware multimodal graph recommendation. SSR follows a four-stage pipeline that decomposes, modulates, fuses and aligns signals within a shared spectral coordinate system. Motivated by the observation that frequency mapping reveals band-wise structure capturing collaborative semantics and modality-specific detail, we organize learning around band-level operations and supervision. Instead of treating bands as static or lightly reweighted features, SSR converts inputs into spectral representations and then reasons about their informativeness and stability. We introduce Spectral Band Masking (SBM), a training-time band-level perturbation with a representation-consistency objective that reduces reliance on brittle bands and improves robustness. We also design a graph-compatible hyperspectral operator (G-HSNO) that models cross-band and cross-modality dependencies through a compact low-rank parameterization. In addition, a spectral contrastive objective promotes band-wise cross-modal consistency without inference overhead. Together, these components provide a cohesive and efficient approach to structured spectral modeling for multimodal graphs.
@@ 2.1 Multi-modal Recommendation
Multimodal recommendation integrates auxiliary content such as images, text and audio to strengthen user preference modeling, especially under sparse supervision He and McAuley (2016); Chen et al. (2019); Yang et al. (2024b, 2025c). Early work augments collaborative filtering with side content Deldjoo et al. (2021); Xu et al. (2023a); Kang et al. (2017). Subsequent methods build modality-aware graph architectures to propagate and fuse signals along user–item relations Wei et al. (2019, 2020); Wang et al. (2021); Yu et al. (2022); Han et al. (2022); Tao et al. (2022). To mitigate cross-modal misalignment and semantic noise, recent approaches employ contrastive learning Zhou et al. (2023b); Yi et al. (2022), regularization Zhang et al. (2021); Zhou and Shen (2023) and intention-aware designs Yang and Yang (2024); Guo et al. (2022). AlignRec Liu et al. (2024b) enforces modality–ID consistency, while PromptMM Wei et al. (2024b) enables lightweight adaptation. Generative strategies, including diffusion models Jiang et al. (2024), adversarial alignment Liang et al. (2022) and large language models Zhao et al. (2025); Xia et al. (2025b); Gu et al. (2025); Xia et al. (2025c), further improve robustness under heterogeneity. In addition, empowered by the reasoning capacity of large language models Chen et al. (2025a); Yang et al. (2025a); Chen et al. (2025b); Ping et al. (2025); Li et al. (2025); Yang et al. (2024a) and multi-agent systems Wang et al. (2024b); Yang and Thomason (2025); Ye et al. (2024a); Chang et al. (2025); Yang et al. (2025b), recent advances such as Rec-GPT4V Liu et al. (2024c) and UniMP Wei et al. (2024a) leverage large vision–language models to strengthen multimodal semantic understanding and improve reasoning over heterogeneous content.
@@ 2.2 Frequency-aware Representation Learning
Frequency-domain modeling has recently emerged as a promising alternative to time- or topology-domain learning in recommendation tasks Fan et al. (2024); Luo et al. (2024c). In sequential modeling, various approaches leverage spectral transformations to capture user intent dynamics and mitigate temporal noise Zhang et al. (2023); Du et al. (2023b), while others propose frequency-aware contrastive objectives or neural filtering for periodic behavior tracking Yang et al. (2025d); Zhang et al. (2025). CFIT4SRec Zhang et al. (2023) adopts frequency-based contrastive learning frameworks, leveraging Fourier transformations to construct robust augmented views and improve representation learning. FEARec Du et al. (2023b) and HMFTRec Fan et al. (2024) add spectral attention and multi-frequency modules to capture low and high frequency signals beyond augmentation. For graph-based recommendation, ChebyCF Kim et al. (2025a), SelfGNN Liu et al. (2024d), and SComGNN Luo et al. (2024a) reformulate GNN propagation as spectral filtering, improving expressivity and denoising capacity. In multimodal settings, SMORE Ong and Khong (2024) explicitly models frequency-specific modality fusion, while FREEDOM Zhou and Shen (2023) and MMGCL Yi et al. (2022) implicitly introduce spectral diversity through augmentation. Nonetheless, most of these methods either treat frequency bands as static views or lack explicit modeling of band reliability and inter-band interactions. We address these gaps by pursuing adaptive, band-aware integration that regulates sensitivity to brittle components and supports interpretable cross-modal fusion.
@@ 3.1 Spectral Transformation of Multimodal Graph Signals
Multimodal recommendation integrates ID embeddings with visual and textual features to learn unified user-item representations. A central challenge is cross-modal heterogeneity, where modalities reside in incompatible spaces and capture semantics at different granularities. Conventional fusion methods such as concatenation or attention operate in the raw feature domain and often ignore the structural dependencies induced by user-item interactions. Viewing these representations as graph signals enables frequency-aware modeling, where low-frequency components capture global preference patterns and high-frequency components encode fine-grained, modality-specific variations. Motivated by this perspective, we transform multimodal features into a unified frequency domain to facilitate band-level alignment and more interpretable fusion, supporting adaptive selection of frequency components for downstream recommendation.
Figure 1: Overall architecture of our proposed framework. The model follows a structured four-stage pipeline: (i) Decomposition performs a modality-specific spectral transform on graph signals to disentangle multi-frequency components; (ii) Modulation applies Spectral Band Masking (SBM) to perturb and down-weight unreliable bands in a task-adaptive manner; (iii) Fusion leverages a low-rank Graph HyperSpectral Neural Operator (G-HSNO) to reason over cross-band and cross-modal dependencies; and (iv) Alignment introduces Spectral Contrastive Regularization (SCR) to enforce semantic consistency and spectral robustness across modalities.
@@
Given a user–item bipartite graph $\mathcal{G}=(\mathcal{N},\mathcal{E})$ with adjacency matrix $\mathbf{A}$ Defferrard et al. (2016); Wu et al. (2019), we use the symmetric normalized Laplacian $\mathbf{L}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$, where $\mathbf{D}$ is the degree matrix. For each node $n\in\mathcal{N}$, the associated signal (e.g., an ID embedding or a multimodal feature) is $\mathbf{x}_{n}\in\mathbb{R}^{d}$. Stacking all nodes yields $\mathbf{X}\in\mathbb{R}^{|\mathcal{N}|\times d}$. We analyze $\mathbf{X}$ in the spectral domain via the eigendecomposition $\mathbf{L}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$, where $\mathbf{U}$ contains orthonormal eigenvectors and $\boldsymbol{\Lambda}$ is diagonal with eigenvalues. The graph Fourier transform (GFT) is $\widehat{\mathbf{X}}=\mathbf{U}^{\top}\mathbf{X}$ and the inverse transform is $\mathbf{X}=\mathbf{U}\widehat{\mathbf{X}}$. Each row of $\widehat{\mathbf{X}}$ corresponds to a graph frequency (eigenmode). In the next step we group these modes into bands for downstream processing. In practice, explicitly computing $\mathbf{U}$ is only feasible for small graphs. We thus adopt a hybrid implementation: for graphs with $|\mathcal{N}|\leq N_{0}$ we compute the full eigendecomposition of $\mathbf{L}$, while for larger graphs we avoid eigen-decomposition and approximate multi-band filtering using Chebyshev polynomials of the scaled Laplacian. In practice, we leverage structural signals from the interaction graph as well as auxiliary user–user and item–item relations to enrich the spectral pipeline.
@@
which yields bands with comparable signal mass and avoids extreme imbalance across modes. Given the orthonormal basis $\mathbf{U}=[\mathbf{u}_{1},\ldots,\mathbf{u}_{|\mathcal{N}|}]$, the reconstruction for band $m$ is $\mathbf{X}^{(m)}=\sum_{k\in\mathcal{F}_{m}}\mathbf{u}_{k}\,\widehat{\mathbf{x}}_{k}^{\top}\,$, where $\mathbf{u}_{k}$ is the $k$-th eigenvector. Each $\mathbf{X}^{(m)}$ captures information at a specific semantic resolution, with lower-index bands concentrating smoother components and higher-index bands emphasizing localized variations. In practice, $M$ is treated as a hyperparameter. When full eigenmodes are available, we partition bands by equal-energy quantiles of the cumulative spectrum. Otherwise, we split the spectrum in the eigenvalue domain and instantiate an $M$-band Chebyshev polynomial filter bank, which provides a scalable approximation to explicit band partitioning without eigendecomposition.
@@
where $\alpha_{n}^{(m)}\in\mathbb{R}_{\geq 0}$ are band-specific weights produced from node information. This allows each node to emphasize informative bands while down-weighting unstable ones. In practice, this aggregation can be implemented either by a lightweight gating module as above or by our subsequent G-HSNO operator, providing a unified and flexible interface for band fusion.